# Generalising Traffic Forecasting to Regions without Traffic Observations

### v1

@@ Introduction
Traffic forecasting is essential for intelligent transportation systems, enabling optimisations such as real-time route planning and transportation scheduling. Accurate traffic forecasting yields substantial social and economic benefits by improving travel efficiency, reducing congestion-related losses, and supporting sustainable urban development. However, high deployment costs of traffic sensors often result in their sparse and limited spatial coverage (Wu et al. 2021), creating a gap between limited traffic observations and the need for fine-grained, wide-coverage forecasting.
@@
To bridge this gap, recent works study kriging and extrapolation. Kriging models estimate current traffic conditions at locations of interest without sensors, i.e., unobserved locations (Zheng et al. 2023; Xu et al. 2025), while extrapolation models take a step further and forecast for such locations (Hu et al. 2023, 2024c). Although these models have produced promising results for scattered unobserved locations (Fig. 1a), they struggle when applied to large continuous regions without traffic sensors (Fig. 1b), i.e., traffic forecasting for unobserved regions (Su et al. 2024b).
@@
Figure 1: Illustration of the problem setting and modelling strategies. Left: (a) Scattered unobserved locations vs. (b) Unobserved region (our focus). Blue bubbles represent observed locations; red hollow bubbles denote unobserved ones; and red solid bubbles denote the target forecasts. Right: Comparison of different modelling strategies: (c) Incorporating static auxiliary features (e.g., learning from locations with similar POIs or geo-coordinates); (d) Integrates physics priors and dynamic external signals (e.g., weather) to improve generalisation (ours).
@@
We consider traffic forecasting for a region without traffic observations that is adjacent to (or enclosed by) observed regions. As Fig. 1b illustrates, the region with red hollow bubbles does not have traffic observations at current time $t$, which is adjacent to the region with observations as denoted by the blue bubbles. We aim to forecast for the unobserved region for a future time (the red bubbles at time $t+1$). This setting is practical due to staged deployment of sensors or unbalanced regional development (Su et al. 2024b).
@@
The state-of-the-art (SOTA) model for this setting, STSM (Su et al. 2024b), masks locations in observed regions that are similar to the unobserved region and is trained to forecast for such locations, with the aim to generalise to unobserved regions. The similarity is defined based on static auxiliary features, i.e., POI categories and geo-coordinates (Fig. 1c), which, however, do not capture the dynamic nature of traffic patterns, thus limiting generalisation capacity.
@@
Another work, KITS (Xu et al. 2025), creates virtual nodes at random locations and trains a model to forecast for such nodes, instead of masking the already-observed locations. It implicitly assumes scattered unobserved locations of a known density (the virtual nodes are created to match this density), thereby introducing a structural prior that limits generalisability to large continuous unobserved regions.
@@
To incorporate physical principles, we exploit the Lighthill–Whitham–Richards (LWR) equation (Lighthill and Whitham 1955; Richards 1956) as a soft learning constraint (i.e., a loss term). LWR governs the relationships between traffic density and flow in a road network. Using it poses two technical challenges: (1) The LWR equation considers both traffic density and flow, while density data are typically unavailable. (2) Applying LWR in model learning requires partial derivative over space, while most traffic forecasting models consider a graph over locations of interest which are discrete and not directly differentiable.
@@
To further utilise dynamic external signals, we exploit global weather observations from ECMWF (Muñoz Sabater 2019), for their universal availability and strong correlation with dynamic traffic patterns (Nigam and Srivastava 2023). We introduce an attention-based fusion module to learn the correlations between weather and traffic patterns.
Beyond introducing external guidance signals, it is also crucial to filter patterns local to individual locations (e.g., induced by a traffic accident) that are non-generalisable to unobserved regions. We propose a spatial grouping module that dynamically learns to group locations based on their intrinsic spatial-temporal patterns. The resulting groups enable GenCast to learn shared patterns of a group and suppress disruptive signals local to individual locations (Lin et al. 2023).
@@
Kriging and extrapolation in traffic forecasting aim to infer current or future observations at unobserved locations (Hu et al. 2023, 2024a, 2024b, 2024c). However, as discussed in the introduction, existing methods often struggle to generalise, particularly when forecasting for regions without traffic observations. Next, we briefly review representative strategies to improve model generalisation.
Physics-guided approaches have been proposed to enhance generalisability of spatial-temporal models (Hwang et al. 2021; Hettige et al. 2024; Verma, Heinonen, and Garg 2024). In the context of traffic forecasting, these approaches fall into two categories: (1) simulating latent dynamics via neural ODEs or energy-based models (Ji et al. 2022; Wang et al. 2022), which typically require fully observed traffic data to initialise hidden model states; and (2) imposing traffic flow constraints (e.g., LWR) via physics-informed neural networks (PINNs) (Shi, Mo, and Di 2021; Zhang et al. 2024), which often assume continuous spatial locations. Our model introduces differentiable spatial embeddings to enable PINNs on traffic graphs which are discrete.
Other studies use external signals to enhance forecasting performance by capturing invariant spatial-temporal patterns from external-domain datasets (Li et al. 2024) or environmental features (e.g., weather (Nigam and Srivastava 2023; Mystakidis and Tjortjis 2024) or events (Su et al. 2024a)). These studies use such signals to help detect irregular events or handle short-term missing data. The use of such signals to guide model generalisation to unobserved regions, combined with advanced spatial-temporal graph networks, remains underexplored. A full discussion of these works is included in Appendix A in the supplementary materials.
@@
Following Su et al. (2024b), we represent a region as a graph $G=(V,E)$, where $V$ denotes $N$ locations of interest and $E$ represents their connections based on spatial proximity. Each location is associated with a feature vector, forming a matrix $\mathbf{L}\in\mathbb{R}^{N\times F}$ that encodes static attributes such as geo-coordinates, road network information, or regional descriptors. The specific features, which can be raw attributes or embeddings, may vary across different methods. For each $v_{i}\in V$, we use $\mathbf{x}_{i}$ to denote the series of traffic observations at location $v_{i}$ and $\mathbf{x}_{i}^{t}\in\mathbb{R}^{C}$ to denote the $C$ different types of observations (e.g., speed and volume) at time $t$, if there are such observations collected.
@@
We collect weather data from the ECMWF ERA5 dataset (Muñoz Sabater 2019), provided in a gridded format with $9\,\text{km}\times 9\,\text{km}$ resolution. For each grid cell $i$ at time $t$, the weather observation is denoted as $\mathbf{x}_{w,i}^{t}\in\mathbb{R}^{C_{w}}$, where $C_{w}=4$ is the number of weather attributes (2-meter temperature, surface net solar radiation, surface runoff, and total precipitation). We denote the full weather observations at time $t$ as $\mathbf{X}_{w}^{t}\in\mathbb{R}^{N_{w}\times C_{w}}$, where $N_{w}$ is the number of grid cells overlapping the input region $G$.
@@
Given region $G=(V,E)=G_{o}+G_{u}$, location features $\mathbf{L}$ corresponding to $V$, traffic observations $\mathbf{X}_{G_{o}}^{t-T+1:t}$ for $G_{o}$ over the past $T$ time steps, and weather observations $\mathbf{X}_{w}^{t-T_{w}+1:t}$ for $G$ over the past $T_{w}$ time steps ($T$ and $T_{w}$ may vary and hence are denoted differently), we aim to learn a function $f$ to forecast the traffic conditions $\hat{\mathbf{X}}_{G_{u}}^{t+1:t+T^{\prime}}$ for $G_{u}$ over the next $T^{\prime}$ time steps:
@@
Fig. 2 presents an overview of GenCast. The backbone of the model (components shown in gray) takes a contrastive learning architecture. At each training epoch, a masked graph view $G_{o}^{m}$ is constructed by randomly masking a subgraph from the observed graph $G_{o}$, yielding two input sequences (i.e., two views): the original $\mathbf{X}_{G_{o}}^{t-T+1:t}$ and the masked $\mathbf{X}_{G_{o}^{m}}^{t-T+1:t}$. Following Su et al. (2024b), we generate pseudo-observations for the masked locations, to enable computing both spatial (i.e., geo-coordinate) proximity-based and temporal (i.e., traffic series) similarity-based adjacency matrices. These allow constructing graph convolution layers (GCNs) to capture spatial dependencies and temporal convolutional networks (TCNs) to capture temporal patterns.
The two views go through the GCNs and TCNs (i.e., a spatial-temporal (ST) model) separately to produce forecasts and graph representations $\mathbf{Z}^{t+T^{\prime}}_{G_{o}}$ and $\mathbf{Z}^{t+T^{\prime}}_{G_{o}^{m}}$, where the representation is taken from the final time step $t+T^{\prime}$ following Liu et al. (2022). A contrastive loss, $L_{cl}$, is applied over $\mathbf{Z}^{t+T^{\prime}}_{G_{o}}$ and $\mathbf{Z}^{t+T^{\prime}}_{G_{o}^{m}}$ to promote consistent forecasts with and without masked (i.e., unobserved) locations, thereby achieving generalisation to the unobserved region. Since both views go through identical pipelines, we omit subscripts $G_{o}$’ and $G^{m}_{o}$’ in the subsequent discussion as long as the context is clear. More details about the backbone contrastive learning process are in Appendix B.
@@
Here, $\theta$ and $\mu$ are hyper-parameters. The contrastive learning loss weighting follows Su et al. (2024b).
@@
LLM-based spatial embedding. SE-L is computed with two steps: (i) location description generation and (ii) embedding generation. Location description generation constructs textual descriptions for each location by prompting an LLM with geo-coordinates, geometric properties of the surrounding area, POI information and attributes of the nearest road segments (see Appendix B). The location description of $S$ tokens, $\mathbf{L}_{t}\in\mathbb{R}^{N_{o}\times S}$, is fed into a frozen LLaMA3 8B Instruct (Meta 2024). The final token embedding from the last hidden layer of the model is SE-L, $\mathbf{L}_{{llm}}\in\mathbb{R}^{N_{o}\times d_{{llm}}}$ ($d_{llm}$ is the embedding dimensionality).
GeoHash-based spatial embeddings. When location features are unavailable beyond the geo-coordinates, we use GeoHash (Niemeyer 2008) to compute spatial embeddings SE-H. There are two main steps: (i) GeoHash string generation and (ii) embedding generation. GeoHash string generation applies GeoHash coding on the geo-coordinates of a location to convert them into a fixed-length alphanumeric string, the length of which decides spatial precision. The embedding generation step then embeds the GeoHash string using a pre-trained character-BERT (Li 2023). The last hidden layer output, discarding the special tokens [CLS] and [SEP], produces $\mathbf{L}^{\prime}_{hash}\in\mathbb{R}^{N_{o}\times S\times d_{bert}}$. Here, $S$ is reused to denote the GeoHash string length, and $d_{{bert}}$ is the embedding dimensionality. To capture semantic dependencies within $\mathbf{L}^{\prime}_{{hash}}$, we feed it into a multi-layer Transformer encoder, and we apply mean pooling along the character dimension of the output to obtain the SE-H, $\mathbf{L}_{{hash}}\in\mathbb{R}^{N_{o}\times d_{{hash}}}$.
@@
To use external weather signals, each traffic location $v_{i}$ is matched to its nearest weather station $s_{j}$ based on proximity:
@@
We adopt spatial grouping to softly cluster locations into latent groups (Fig. 3), enabling GenCast to capture shared patterns of a group and filtered out ad hoc patterns at individual locations, with the help of an entropy regularisation term.
@@
$\mathbf{W}_{s}=\text{Softmax}\left(-\text{cdist}(\mathbf{Z}^{l},\mathbf{C})\right)\in\mathbb{R}^{(N\cdot cg)\times(sg\cdot cg)}.$ (5)
@@
The physics-informed module introduces a constraint with the LWR model (Lighthill and Whitham 1955; Richards 1956). LWR describes the evolution of traffic density over space (location $l$) and time ($t$) using a conservation law formulated as Eq. 7, where $\rho=\rho(l,t)$ denotes traffic density, andx=x(l,t)x=x(l,t) represents traffic velocity (i.e., speed).
@@
As traffic density observations are not commonly available, we rewrite the equation using velocity, assuming a closed system with a functional density-velocity relationship (Greenshields 1935; Xiong, Zhou, and Bennett 2023): $x=x_{fspd}\left(1-\frac{\rho}{\rho_{{max}}}\right)$, where $x_{fpsd}$ denotes the free flow speed – the speed at which vehicles travel under low traffic density, and $\rho_{max}$ denotes the maximum traffic density, where vehicles are fully packed. This can also be written as:
@@
To guide GenCast to follow the physical principles, we use the Huber loss (Huber 1992) to minimise the physics residual (i.e., to penalise violations of the physical law):
$\mathcal{L}_{{phy}}=\mathrm{Huber}(R,\delta)$ (11)
@@
where $\hat{\mathbf{X}}=\hat{\mathbf{X}}_{G_{o}^{m}}^{t+1:t+T^{\prime}}$ denotes model forecasts based on the masked graph, and $\mathbf{X}_{{fspd}}\in\mathbb{R}^{N\times 1}$ represents estimated free-flow speed at each location (Xiong, Zhou, and Bennett 2023). The partial derivatives with respect to $\mathbf{TE}_{{enc}}$ and $\mathbf{L}_{{enc}}$ are computed via automatic differentiation.
@@
We evaluate GenCast on four highway datasets (PEMS-Bay, PEMS07, PEMS08, and METR-LA) and one urban dataset (Melbourne), with 5-min or 15-min intervals. Dataset details, including statistics, processing, and visualisations, are in Appendix C.1. We use ERA5-Land weather data (Muñoz Sabater 2019) (hourly, $\times9km$) with four traffic-related variables: temperature, solar radiation, precipitation, and runoff. Region and road network data for LLM-based embeddings are from OpenStreetMap (2018).
Following prior work (Zheng et al. 2023; Su et al. 2024b), we use traffic records from the past two hours to forecast for the next two hours, i.e., $T=T^{\prime}=2$ hours. Each dataset is split into training, validation, and test sets in a 4:1:5 spatial ratio, ensuring spatial adjacency within each split. Locations in the training and validation sets are treated as observed, while test locations are unobserved. The space-based split is performed horizontally or vertically based on geo-coordinates. We generate four spatial splits per dataset and report results on average. Temporally, the first 70% of data is used for training, and the remaining 30% for testing.
@@
We compare with a transductive Kriging model GE-GAN (Xu et al. 2020), inductive Kriging models IGNNK (Wu et al. 2021), INCREASE (Zheng et al. 2023) and KITS (Xu et al. 2025), and the SOTA model for unobserved region forecasting STSM (Su et al. 2024b).
@@
[…]
Table 1: Overall model performance. $\downarrow$” (and $\uparrow$”) indicates that lower (and larger) values are better. The best baseline results are underlined, and the best results are in bold. “Improve” means the errors reduced by GenCast-L compared with the best baseline model.

### v2

@@ Introduction
Traffic forecasting is essential for intelligent transportation systems, enabling optimisations such as real-time route planning and transportation scheduling. Accurate traffic forecasting yields substantial social and economic benefits by improving travel efficiency, reducing congestion-related losses, and supporting sustainable urban development (gman; gamcn). However, high deployment costs of traffic sensors often result in their sparse and limited spatial coverage (ignnk), creating a gap between limited traffic observations and the need for fine-grained, wide-coverage forecasting.
@@
To bridge this gap, recent works study kriging and extrapolation. Kriging models estimate current traffic conditions at locations of interest without sensors, i.e., unobserved locations (increase; kits), while extrapolation models take a step further and forecast for such locations (STGNP; stgp). Although these models have produced promising results for scattered unobserved locations (Fig. 1a), they struggle when applied to large continuous regions without traffic sensors (Fig. 1b), i.e., traffic forecasting for unobserved regions (stsm).
@@
Figure 1: Illustration of the problem setting and modelling strategies. Upper: (a) Scattered unobserved locations vs. (b) Unobserved region (our focus). Blue bubbles represent observed locations; red hollow bubbles denote unobserved ones; and red solid bubbles denote the target forecasts. Lower: Comparison of different modelling strategies: (c) Incorporating static auxiliary features (e.g., learning from locations with similar POIs or geo-coordinates); (d) Integrates physics priors and dynamic external signals (e.g., weather) to improve generalisation (ours).
@@
We consider traffic forecasting for a region without traffic observations that is adjacent to (or enclosed by) observed regions. As Fig. 1b illustrates, the region with red hollow bubbles does not have traffic observations at current time $t$, which is adjacent to the region with observations as denoted by the blue bubbles. We aim to forecast for the unobserved region for a future time (the red bubbles at time $t+1$). This setting is practical due to the staged deployment of sensors or unbalanced regional development (stsm).
@@
The state-of-the-art (SOTA) model for this setting, STSM (stsm), masks locations in observed regions that are similar to the unobserved region and is trained to forecast for such locations, with the aim to generalise to unobserved regions. The similarity is defined based on static auxiliary features, i.e., POI categories and geo-coordinates (Fig. 1c), which, however, do not capture the dynamic nature of traffic patterns, thus limiting generalisation capacity.
@@
Another work, KITS (kits), creates virtual nodes at random locations and trains a model to forecast for such nodes, instead of masking the already-observed locations. It implicitly assumes scattered unobserved locations of a known density (the virtual nodes are created to match this density), thereby introducing a structural prior that limits generalisability to large continuous unobserved regions.
@@
To incorporate physical principles, we exploit the Lighthill–Whitham–Richards (LWR) equation (lwr1; lwr2) as a soft learning constraint (i.e., a loss term). LWR governs the relationships between traffic density and flow in a road network. Using it poses two technical challenges: (1) The LWR equation considers both traffic density and flow, while density data are typically unavailable. (2) Applying LWR in model learning requires partial derivative over space, while most traffic forecasting models consider a graph over locations of interest which are discrete and not directly differentiable.
@@
To further utilise dynamic external signals, we exploit global weather observations from ECMWF (ecmwf_era5), for their universal availability and strong correlation with dynamic traffic patterns (nigam2023hybrid). We introduce an attention-based fusion module to learn the correlations between weather and traffic patterns.
Beyond introducing external guidance signals, it is also crucial to filter patterns local to individual locations (e.g., induced by a traffic accident) that are non-generalisable to unobserved regions. We propose a spatial grouping module that dynamically learns to group locations based on their intrinsic spatial-temporal patterns. The resulting groups enable GenCast to learn shared patterns of a group and suppress disruptive signals local to individual locations (lin2023diversifying).
@@
Kriging and extrapolation in traffic forecasting aim to infer current or future observations at unobserved locations (STGNP; dualSTN; ustd; stgp). However, as discussed in the introduction, existing methods often struggle to generalise, particularly when forecasting for regions without traffic observations. Next, we briefly review representative strategies to improve model generalisation.
Physics-guided approaches have been proposed to enhance generalisability of spatial-temporal models (hwang2021climate; hettigeairphynet; ClimODE). In the context of traffic forecasting, these approaches fall into two categories: (1) simulating latent dynamics via neural ODEs or energy-based models (STDEN; st-pef), which typically require fully observed traffic data to initialise hidden model states; and (2) imposing traffic flow constraints (e.g., LWR) via physics-informed neural networks (PINNs) (raissi2019physics; shi2021physics; zhang2024physics), which often assume continuous spatial locations. Our model introduces differentiable spatial embeddings to enable PINNs on traffic graphs which are discrete.
Other studies use external signals to enhance forecasting performance by capturing invariant spatial-temporal patterns from external-domain datasets (li2024opencity) or environmental features (e.g., weather (mystakidis2024traffic) or events (zhou2024sdformer; dualcast; ruan2025retrieval)). These studies use such signals to help detect irregular events or handle short-term missing data. The use of such signals to guide model generalisation to unobserved regions, combined with advanced spatial-temporal graph networks, remains underexplored. A full discussion of these works is included in Appendix A.
@@
Following stsm, we represent a region as a graph $G=(V,E)$, where $V$ denotes $N$ locations of interest and $E$ represents their connections based on spatial proximity. Each location is associated with a feature vector, forming a matrix $\mathbf{L}\in\mathbb{R}^{N\times F}$ that encodes static attributes such as geo-coordinates, road network information, or regional descriptors. The specific features, which can be raw attributes or embeddings, may vary across different methods. For each $v_{i}\in V$, we use $\mathbf{x}_{i}$ to denote the series of traffic observations at location $v_{i}$ and $\mathbf{x}_{i}^{t}\in\mathbb{R}^{C}$ to denote the $C$ different types of observations (e.g., speed and volume) at time $t$, if there are such observations collected.
@@
We collect weather data from the ECMWF ERA5 dataset (ecmwf_era5), provided in a gridded format with $9\,\text{km}\times 9\,\text{km}$ resolution. For each grid cell $i$ at time $t$, the weather observation is denoted as $\mathbf{x}_{w,i}^{t}\in\mathbb{R}^{C_{w}}$, where $C_{w}=4$ is the number of weather attributes (2-meter temperature, surface net solar radiation, surface runoff, and total precipitation). We denote the full weather observations at time $t$ as $\mathbf{X}_{w}^{t}\in\mathbb{R}^{N_{w}\times C_{w}}$, where $N_{w}$ is the number of grid cells overlapping the input region $G$.
@@
Given region $G=(V,E)=G_{o}\cup G_{u}$, location features $\mathbf{L}$ corresponding to $V$, traffic observations $\mathbf{X}_{G_{o}}^{t-T+1:t}$ for $G_{o}$ over the past $T$ time steps, and weather observations $\mathbf{X}_{w}^{t-T_{w}+1:t}$ for $G$ over the past $T_{w}$ time steps ($T$ and $T_{w}$ may vary and hence are denoted differently), we aim to learn a function $f$ to forecast the traffic conditions $\hat{\mathbf{X}}_{G_{u}}^{t+1:t+T^{\prime}}$ for $G_{u}$ over the next $T^{\prime}$ time steps:
@@
Fig. 2 presents an overview of GenCast. The backbone (in gray) adopts a contrastive architecture with a pseudo-observation generator, following (stsm). Each epoch constructs a masked graph $G_{o}^{m}$ by randomly masking a subgraph of $G_{o}$, producing two input views: original $\mathbf{X}{G_{o}}^{t-T+1:t}$ and masked $\mathbf{X}{G_{o}^{m}}^{t-T+1:t}$. Pseudo-observations are generated for masked locations to compute spatial (geo-coordinate) and temporal (traffic-series) similarity matrices, which feed GCNs and TCNs to capture respective dependencies.
The two views go through the GCNs and TCNs (i.e., a spatial-temporal (ST) model) separately to produce forecasts and graph representations $\mathbf{Z}^{t+T^{\prime}}_{G_{o}}$ and $\mathbf{Z}^{t+T^{\prime}}_{G_{o}^{m}}$, where the representation is taken from the final time step $t+T^{\prime}$ following whdocl. A contrastive loss, $L_{cl}$, is applied over $\mathbf{Z}^{t+T^{\prime}}_{G_{o}}$ and $\mathbf{Z}^{t+T^{\prime}}_{G_{o}^{m}}$ to promote consistent forecasts with and without masked (i.e., unobserved) locations, thereby achieving generalisation to the unobserved region. Since both views go through identical pipelines, we omit subscripts $G_{o}$’ and $G^{m}_{o}$’ in the subsequent discussion as long as the context is clear. More details about the backbone contrastive learning process are in Appendix B.
@@
Here, $\theta$ and $\mu$ are hyper-parameters. The contrastive learning loss weighting follows stsm.
@@
LLM-based spatial embedding. SE-L is computed with two steps: (i) location description generation and (ii) embedding generation following (sun2025flexireg). Location description generation constructs textual descriptions for each location by prompting an LLM with geo-coordinates, geometric properties of the surrounding area, POI information and attributes of the nearest road segments (see Appendix B). The location description of $S$ tokens, $\mathbf{L}_{t}\in\mathbb{R}^{N_{o}\times S}$, is fed into a frozen LLaMA3 8B Instruct (llama2024). The final token embedding from the last hidden layer of the model is SE-L, $\mathbf{L}_{{llm}}\in\mathbb{R}^{N_{o}\times d_{{llm}}}$ ($d_{llm}$ is the embedding dimensionality).
GeoHash-based spatial embeddings. When location features are unavailable beyond the geo-coordinates, we use GeoHash (geohash) to compute spatial embeddings SE-H. There are two main steps: (i) GeoHash string generation and (ii) embedding generation. GeoHash string generation applies GeoHash coding on the geo-coordinates of a location to convert them into a fixed-length alphanumeric string, the length of which decides spatial precision. The embedding generation step then embeds the GeoHash string using a pre-trained character-BERT (lhy_charbert_2023). The last hidden layer output, discarding the special tokens [CLS] and [SEP], produces $\mathbf{L}^{\prime}_{hash}\in\mathbb{R}^{N_{o}\times S\times d_{bert}}$. Here, $S$ is reused to denote the GeoHash string length, and $d_{{bert}}$ is the embedding dimensionality. To capture semantic dependencies within $\mathbf{L}^{\prime}_{{hash}}$, we feed it into a multi-layer Transformer encoder, and we apply mean pooling along the character dimension of the output to obtain the SE-H, $\mathbf{L}_{{hash}}\in\mathbb{R}^{N_{o}\times d_{{hash}}}$.
@@
For external weather signals, each traffic location $v_{i}$ is matched to its nearest weather station $s_{j}$ based on proximity:
@@
We adopt spatial grouping to softly cluster locations into latent groups (Fig. 9), enabling GenCast to capture shared group-level patterns and filter out ad hoc patterns at individual locations, with the help of an entropy regularisation term.
@@
$\mathbf{W}_{s}=\text{Softmax}\left(-\text{cdist}(\mathbf{Z}^{l},\mathbf{C})\right)\in\mathbb{R}^{(N_{o}\cdot cg)\times(sg\cdot cg)}.$ (5)
@@
The physics-informed module introduces a constraint with the LWR model (lwr1; lwr2). LWR describes the evolution of traffic density over space (location $l$) and time ($t$) using a conservation law formulated as Eq. 7, where $\rho=\rho(l,t)$ denotes traffic density, andx=x(l,t)x=x(l,t) represents traffic velocity (i.e., speed).
@@
As traffic density observations are not commonly available, we rewrite the equation using velocity, assuming a closed system with a functional density-velocity relationship (greenshields; fspd): $x=x_{fspd}\left(1-\frac{\rho}{\rho_{{max}}}\right)$, where $x_{fpsd}$ denotes the free flow speed – the speed at which vehicles travel under low traffic density, and $\rho_{max}$ denotes the maximum traffic density, where vehicles are fully packed. This can also be written as:
@@
As the system is not assumed to be closed, the model enforces local flow conservation across connected segments consistent with LWR principles. The Huber loss (huber1992robust) penalises violations of flow conservation to enforce these physical constraints
$\mathcal{L}_{{phy}}=\mathrm{Huber}(R,\delta).$ (11)
@@
where $\hat{\mathbf{X}}=\hat{\mathbf{X}}_{G_{o}^{m}}^{t+1:t+T^{\prime}}$ denotes model forecasts based on the masked graph, and $\mathbf{X}_{{fspd}}\in\mathbb{R}^{N\times 1}$ represents estimated free-flow speed at each location (fspd). The partial derivatives with respect to $\mathbf{TE}_{{enc}}$ and $\mathbf{L}_{{enc}}$ are computed via automatic differentiation.
@@
We evaluate GenCast on four highway datasets (PEMS-Bay, PEMS07, PEMS08, and METR-LA) and one urban dataset (Melbourne), with 5-min or 15-min intervals. Dataset details, including statistics, processing, and visualisations, are in Appendix C.1. We use ERA5-Land weather data (ecmwf_era5) (hourly, $\times9km$) with four traffic-related variables: temperature, solar radiation, precipitation, and runoff. Region and road network data for LLM-based embeddings are from osm.
Following prior work (increase; stsm), we use traffic records from the past two hours to forecast for the next two hours, i.e., $T=T^{\prime}=2$ hours. Each dataset is split into training, validation, and test sets in a 4:1:5 spatial ratio, ensuring spatial adjacency within each split. Locations in the training and validation sets are treated as observed, while test locations are unobserved. The space-based split is performed horizontally or vertically based on geo-coordinates. We generate four spatial splits per dataset and report results on average. Temporally, the first 70% of data is used for training, and the remaining 30% for testing.
@@
We compare with a transductive Kriging model GE-GAN (ge-gan), inductive Kriging models IGNNK (ignnk), INCREASE (increase) and KITS (kits), and the SOTA model for unobserved region forecasting STSM (stsm).
@@
    Dataset Metric GE-GAN IGNNK INCREASE STSM KITS GenCast-H GenCast-L Improve
@@
Table 1: Overall model performance. $\downarrow$$\uparrow$” indicates lower/larger is better. The best baseline results are underlined, and the best results are in bold. “Improve” means the errors reduced by GenCast-L compared with the best baseline model.